OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x). Table: x f(x) 0 30 1 32 2 34 Equation: g(x) g(x)=8x-25 Part 1. Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). Part 2. Which function has a great y-intercept? Please help!

OpenStudy (campbell_st):

ok... did you get an equation from the table..?

OpenStudy (anonymous):

Not yet. I'm not sure how to. :/

OpenStudy (campbell_st):

well the equation for the table is y = 2x + 30 slope (32 -30)/(1-0) = 2 when x = 0, y = 30 so now you have the equation what can you say able the slopes..?

OpenStudy (anonymous):

I'm not sure. I don't understand very well.

OpenStudy (campbell_st):

ok... for any linear equation the slope intercept form is y = mx + b m is the slope of the line b is the y-intercept, or where the line cuts the y axis so f(x) has a slope of 2 g(x) has a slope of 8 the larger the number for the slope, the steeper the line.  so the questions... so to answer question 1. you can say, the point slope form of a line is y = mx + b. f(x) has a slope of 2 and g(x) has a slope of 8, g(x) is the line with the steepest slope. Question 2 f(x) has a y intercept of 30 and g(x) has a y intercept of -25... so f(x) has the greater intercept. hope this helps

OpenStudy (anonymous):

Thanks, this help A LOT! :D